in a simple lottery system in which three one-digit numbers are randomly selected what is the probability of getting the first two numbers right and the last number wrong? (numbers can be repeated and 0 is allowed)
Answer #1

total number of ways 3 digit can be selected =10*10*10=1000

number of ways of getting  first two numbers right and the last number wrong =1*1*9=9

therefore probability of getting the first two numbers right and the last number wrong =9/1000=0.009